x = -4y
3x + 2y = 20

Answer by jim_thompson5910(35256)

Hint:

3x + 2y = 20

3(-4y) + 2y = 20 ... plug in x = -4y (ie replace 'x' with '-4y')

Now solve 3(-4y) + 2y = 20 for y

Once you have the solution for y, use it to find the solution for x.